ABSTRACT. This is a mainly expository paper concerning two different definitions of the angle between a pair of subspaces of a Hilbert space, certain basic results which hold for these angles, and a few of the many applications of these notions. The latter include the rate of convergence of the method of cyclic projections, existence and uniqueness of abstract splines, and the product of operators with closed range.